40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8749     Entry time: Tue Jun 25 23:49:04 2013
Author: Annalisa 
Type: Update 
Category: SUS 
Subject: ETMY Oplev 

I had some problem with the Oplev Servo today. I was working at the mode matching fine tuning and I left the Oplev servo enabled while aligning.

When I opened the Yend table lids, the Oplev beam started moving on the QPD and the Oplev servo didn't help in stopping the mirror movement, but it increased it.

So, the mirror was oscillating at a frequency of a few Hz

Koji suggested that maybe the shaking is due to the air conditioning moving the beam, so the servo tries to feed back the signal to the mirror, even if the mirror doesn't actually move.

I also measured the transfer function for the Oplev, but it didn't show any strange behavior.

ELOG V3.1.3-